Please help me solve this problem: Three times the sum of a number and 3 is the same as three times the number divided by 3
Answer by Boreal(15235)

x=number
3(x+3)=3(x/3)
3x+9=x
2x=-9
x=-4.5 ANSWER
3(-1.5)=-4.5/3